OpenStudy (lgbasallote):

a 3.00 kg object is moving in a plane, with its x and y coordinates given by x = 5t^2 - 1 and y = 3t^2 +2, where x and y are in meters and t is in seconds. find the magnitude of the net force acting on this object at t = 2.00 s do i just substitute t?

OpenStudy (anonymous):

is the answer \[3 \sqrt{136} ?\]

OpenStudy (anonymous):

thats right

OpenStudy (lgbasallote):

why are you asking me the answer o.O

OpenStudy (lgbasallote):

i do not know how to do it and im asked the answer..that's bizarre lol

OpenStudy (anonymous):

I thought u did have the answer! most questions give the final answer, that's why.

OpenStudy (lgbasallote):

mehh not this one..

OpenStudy (anonymous):

anyway, u don't need to substitute t at all. acceleration comes out to be a constant. multiply that with mass and u'll get the answer.

OpenStudy (lgbasallote):

how do i find acceleration? 2d/t^2?

OpenStudy (anonymous):

that's right. so u'll get the x and y components of accel.

OpenStudy (lgbasallote):

hmm i cant visualize it...would you show a demo?

OpenStudy (anonymous):

how abt u think of it as a projectile motion. we have constant acceleration. and displacement and velocity are also in 2D - x and y. so, instead of directly giving the values of disp and vel, the values of their components are given in terms of t. does this help?

OpenStudy (lgbasallote):

ohhhhh definitely!!

OpenStudy (anonymous):

so u got the sum?

OpenStudy (lgbasallote):

so i will also have Fx and Fy?

OpenStudy (anonymous):

yeah....u can go that way. and then Fnet = sqrt( Fx^2 + Fy^2)

OpenStudy (lgbasallote):

oh of course

OpenStudy (anonymous):

or else, when u ax and ay....calculate the net accel using the same formula.

OpenStudy (lgbasallote):

wait...does t = 2 s matter? i differentiated x and i got a constant acceleration

OpenStudy (anonymous):

*find

OpenStudy (lgbasallote):

hm?

OpenStudy (anonymous):

that's what i said b4. it doesn't matter.

OpenStudy (lgbasallote):

yeah..just making sure so Fx = 30 N Fy = 18 N so \[\sum F = \sqrt{8^2 + 30^2}?\]

OpenStudy (anonymous):

yup. that's right. just that, under the square root, it should be 18^2...not 8^2

OpenStudy (lgbasallote):

oh lol of course thanks :D

OpenStudy (anonymous):

np :)